OpenStudy (anonymous):

which ratio is proportional to 18/21

OpenStudy (anonymous):

6/7

OpenStudy (anonymous):

correct ! 6/7

OpenStudy (anonymous):

Just do this : 

OpenStudy (anonymous):

Ore : 18/21=36/42

OpenStudy (anonymous):

its ether A.1/3 B.36/48 C.2/3 or D.30/35

OpenStudy (anonymous):

no @E.ali i think we need the shortest form

OpenStudy (anonymous):

so D is correct

OpenStudy (anonymous):

No ! there sno diffrence !:)

OpenStudy (anonymous):

30/35 = 6/7

OpenStudy (anonymous):

What s the D ???

OpenStudy (anonymous):

the MC!

OpenStudy (anonymous):

What ???

OpenStudy (anonymous):

Multiply Choises

OpenStudy (anonymous):

12/15= 8/x NEW QUESTION A.9 B. 10 C.12 D.16

OpenStudy (anonymous):

12/15= 4/5 and we have : 4/5=8/x so x=10 Got it ??:)